OpenStudy (anonymous):

2(x+a)=4b solve for a

OpenStudy (texaschic101):

2(x + a) = 4b --- distribute 2x + 2a = 4b -- subtract 2x 2a = 4b - 2x -- divide by 2 a = 2b - x